# LMM-PCQA: Assisting Point Cloud Quality Assessment with LMM

### v1

@@ Abstract.
Although large multi-modality models (LMMs) have seen extensive exploration and application in various quality assessment studies, their integration into Point Cloud Quality Assessment (PCQA) remains unexplored. Given LMMs’ exceptional performance and robustness in low-level vision and quality assessment tasks, this study aims to investigate the feasibility of imparting PCQA knowledge to LMMs through text supervision. To achieve this, we transform quality labels into textual descriptions during the fine-tuning phase, enabling LMMs to derive quality rating logits from 2D projections of point clouds. To compensate for the loss of perception in the 3D domain, structural features are extracted as well. These quality logits and structural features are then combined and regressed into quality scores. Our experimental results affirm the effectiveness of our approach, showcasing a novel integration of LMMs into PCQA that enhances model understanding and assessment accuracy. We hope our contributions can inspire subsequent investigations into the fusion of LMMs with PCQA, fostering advancements in 3D visual quality analysis and beyond.
@@
copyright: acmcopyrightjournalyear: 2018doi: 10.1145/1122445.1122456conference: *; Arxiv Version; For Reviewbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Visualization design and evaluation methodsccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Point clouds are increasingly used across diverse real-world scenarios, including virtual/augmented reality (Park et al., 2008; Liu et al., 2020; Gu et al., 2019), autonomous vehicles (Cui et al., 2021), and video post-production (Mekuria et al., 2016). This surge is attributed to their adeptness in three-dimensional representation. Consequently, significant research efforts have been channeled towards enhancing high-level areas like point cloud classification (Grilli et al., 2017; Ku et al., 2018; Wang and Jia, 2019; Vora et al., 2020; Xie et al., 2020; Yoo et al., 2020; Chen et al., 2020), detection (Cui et al., 2021), and segmentation (Cheng et al., 2021; Liu et al., 2022). Meanwhile, as a key component for ensuring the point cloud quality, point cloud quality assessment (PCQA) has seen comparable advancements as well during the last decade. PCQA’s objective is to appraise point clouds’ visual quality, a pivotal factor for refining simplification and compression strategies in practical applications (Fan et al., 2022), and to elevate the Quality of Experience (QoE) for end-users. The complexity of PCQA is further compounded when assessing point clouds depicting realistic objects or individuals, due to their intricate geometric structures and the dense aggregation of points, often augmented with color attributes, posing additional challenges. Generally, PCQA methods are divided into three categories, depending on their reliance on reference point clouds: Full-Reference PCQA (FR-PCQA), Reduced-Reference PCQA (RR-PCQA), and No-Reference PCQA (NR-PCQA) respective. The availability of pristine reference point clouds is often limited in real-world applications, emphasizing the demand for NR-PCQA approaches, which is why our research primarily focuses on NR-PCQA.
@@
Secondly, to address the potential insensitivity to geometric distortions (e.g., compression, downsampling) when only analyzing projections, we propose the extraction of multi-scale structural features. This approach enhances the LMM-PCQA’s holistic comprehension of point cloud visual quality. The point clouds are converted into quality-aware structural domains, a technique validated for effective quality feature extraction in prior research (Alexiou et al., 2021; Zhou et al., 2023; Zhang et al., 2022). We modify the scale parameters in the k-nearest neighbors (k-NN) algorithm to offer a multi-scale perspective, aligning with the human vision system’s perception mechanism. Subsequently, key statistical parameters are used to quantify structural distortions within these domains. Finally, we combine the LMM’s evaluative results and the structural features, utilizing support vector regression (SVR) to derive the quality values. The experimental outcomes affirm that our LMM-PCQA model is on par with, or surpasses, current leading PCQA methods.
@@ 2.1. LMM for Quality Assessment
Recent studies have begun to explore the utilization of LMMs for visual quality assessment, marking a significant shift in the field. For instance, X-IQE (Chen, 2023) employs LMMs to assess the quality of text-to-image generation methods, leveraging a chain of thoughts strategy to understand and evaluate visual outputs. Following this, the Q-Bench study (Wu et al., 2024a) presents a binary softmax approach that enables LMMs to compute quantifiable quality scores. This is achieved by utilizing softmax pooling on the logits corresponding to the tokens good and poor, facilitating a more nuanced assessment of visual quality. Further refining this methodology, Q-Instruct (Wu et al., 2024b) demonstrates the potential of fine-tuning LMMs using a text-based question-answering format, targeting specific low-level visual queries to improve their accuracy in visual quality evaluation. Drawing inspiration from these developments, the Q-Align (Wu et al., 2023) framework mimics human evaluation mechanisms and post-assessment recalibrations in visual quality scoring, thereby attaining exceptional performance metrics. This progression in research underscores the evolving capabilities of LMMs in the nuanced domain of visual quality assessment, setting a new benchmark for future explorations. These works have fully illustrated the potential of LMMs’ quality assessment abilities on 2D media such as images and videos. However, no effort has been put into adapting LMMs to the 3D quality assessment field.
@@ 2.2. PCQA Development
In the initial phase of PCQA exploration, the MPEG group introduces key Full-Reference PCQA (FR-PCQA) methods such as p2point (Mekuria et al., 2016) and p2plane (Tian et al., 2017), aimed at evaluating point cloud quality. They later present a point-based PSNR-yuv technique to handle colored point clouds (Torlig et al., 2018). However, these early attempts face some critical challenges in accurately capturing complex distortions through point-level differences. To overcome these challenges mentioned above, more sophisticated FR-PCQA metrics emerge, including PCQM (Meynet et al., 2020), GraphSIM (Yang et al., 2020) and PointSSIM (Alexiou and Ebrahimi, 2020), which integrate structural features to significantly improve performance.
Inspired by recent advances in no-reference image and video quality assessment (NR-I/VQA) (Zhang et al., 2015; Liu et al., 2015; Gu et al., 2017; Zhang and Liu, 2022; Zhang et al., 2022, 2021), and to cater to a broader range of real-world applications, various NR-PCQA methods have been carried out during the last decade. Chetouani $et$ $al.$ (Chetouani et al., 2021) apply classical Convolutional Neural Network (CNN) models for quality regression with the assistance of patch-wise handcrafted features. PQA-net (Liu et al., 2021) adopts multi-view projection for feature extraction, while Zhang $et$ $al.$ (Zhang et al., 2022) predict quality-aware parameters by analyzing distributions of geometry and color attributes. Furthermore, Liu $et$ $al.$ (Liu et al., 2022) utilize an end-to-end sparse CNN for direct quality prediction, and Zhou $et$ $al.$ (Zhou et al., 2022) focus on structure-guided resampling for extracting relevant features. Some researchers attempt to convert point clouds into videos to employ video quality assessment (VQA) techniques for evaluating perceptual quality (Fan et al., 2022; Zhang et al., 2022). To mitigate computational complexity, GMS-3DQA approach (Zhang et al., 2023) integrates multi-projections into a singular quality map to facilitate feature extraction. Yang $et$ $al.$ (Yang et al., 2022) extend quality assessment methodologies from natural images to point cloud rendering images via domain adaptation techniques. Additionally, Wang $et$ $al.$ (Wang et al., 2023) employs a sophisticated non-local geometry and color gradient aggregation graph model for accurate quality estimation. Furthermore, to provide a more comprehensive learning framework for point clouds from both color and geometry perspectives, MM-PCQA (Zhang et al., 2023) and pmBQA (Xie et al., 2023) suggest leveraging a multi-modal learning approach to improve PCQA outcomes. In the evolution of PCQA, the potential of text modality for quality representation remains largely untapped. Investigating the approach of leveraging text supervision to facilitate model quality assessment learning emerges as an innovative strategy.
@@ 3.1.1. Projection Acquisition
Consider a colored point cloud $\mathbf{P}={(p_{i},c_{i})}_{i=1}^{N}$, where $p_{i}\in\mathbb{R}^{1\times 3}$ represents the single point consisting of geometry coordinates, $c_{i}\in\mathbb{R}^{1\times 3}$ represents the RGB color attributes, and $N$ denotes the total count of points. We then adopt the conventional cube-like viewpoints configuration, widely used in the standard point cloud compression scheme MPEG VPCC (Graziosi et al., 2020). As illustrated in Fig. 2, six orthogonal viewpoints are utilized, each mapping to one of the cube’s six faces for generating the projections. For a point cloud $\mathbf{P}$, the rendering process is derived as:
@@ 3.1.2. How to inject PCQA knowledge into LMMs ?
LMM has shown competitive performance in 2D image/video quality assessment (Wu et al., 2023, 2024a, 2024b), therefore it is feasible to apply LMM for PCQA tasks by taking the point cloud as a sequence of projections. Then it is vital to solve the core problem of PCQA Knowledge Injection. Following the common approaches of training LMMs (Ye et al., 2023; Liu et al., 2023), it is natural to come up with the solution of instructing the LMM with question-answer pairs regarding PCQA issues. Thus we carry out the specific prompt structure as follows:
@@ 3.2. Structual Feature Extraction
The inadequacy of single-modality information from projections for point cloud quality evaluation has been established (Zhang et al., 2023; Xie et al., 2023). Therefore, our approach enhances the accuracy of LMM evaluation results (projections only) by integrating geometric structural features, aiming for a more detailed and accurate assessment.
@@ 3.2.1. Structural Domain
where $Lin(p_{i})$ and $Pla(p_{i})$ represent the linearity and planarity values for point $p_{i}$. The chosen structural domains (linearity, planarity) have been demonstrated to exhibit a strong correlation with geometric visual losses, such as compression and downsampling, and have been extensively utilized in numerous PCQA tasks (Alexiou et al., 2021; Zhou et al., 2023; Zhang et al., 2022).
@@ 3.2.2. Multi-scale Perception
The multi-scale nature of point cloud visual perception has been noted in the literature (Zhang et al., 2021). To account for this, we compute structural domains across various scales by varying the scale parameter $k$ in the $\boldsymbol{\rm{KNN}}(\cdot)$ process. This approach allows us to derive the multi-scale structural domains as follows:
@@ 4.1. Validation Databases
To assess the efficacy of the proposed method, we employ the SJTU-PCQA database (Yang et al., 2020), the Waterloo point cloud assessment database (WPC)(Liu et al., 2022), and the WPC2.0 database (Liu et al., 2021) for validation. The SJTU-PCQA database contains 9 reference point clouds, subjected to seven distortion types (compression, color noise, geometric shift, down-sampling, and three mixed distortions) at six levels, yielding a total of 378 ($9\times 7\times 6$) distorted point clouds. The WPC database includes 20 reference point clouds, each modified by four distortions (down-sampling, Gaussian white noise, Geometry-based Point Cloud Compression (G-PCC), and Video-based Point Cloud Compression (V-PCC)), resulting in 740 ($20\times 37$) distorted point clouds. Meanwhile, the WPC2.0 database features 16 reference point clouds, each undergoing 25 different V-PCC degradation settings, leading to 400 ($16\times 25$) distorted point clouds.
@@ 4.2.2. Validation Strategy
Following the methodologies in (Fan et al., 2022; Zhang et al., 2023; Chai et al., 2024), we utilize a k-fold cross-validation approach in our experiments to ensure a dependable performance evaluation of our proposed method. The SJTU-PCQA, WPC, and WPC2.0 databases consist of 9, 20, and 16 point cloud groups, respectively, leading us to adopt 9-fold, 5-fold, and 4-fold cross-validation for these databases to achieve an approximate 8:2 train-test split. The average of the performance metrics is considered the definitive result. It is crucial to note that the training and testing sets are mutually exclusive to prevent content overlap. For FR-PCQA methods, which do not require training, we evaluate them using the same test sets and report the average performance.
@@ 4.3. Competitors
  • The FR-PCQA methods include MSE-p2point (MSE-p2po) (Mekuria et al., 2016), Hausdorff-p2point (HD-p2po) (Mekuria et al., 2016), MSE-p2plane (MSE-p2pl) (Tian et al., 2017), Hausdorff-p2plane (HD-p2pl) (Tian et al., 2017), PSNR-yuv (Torlig et al., 2018), PCQM (Meynet et al., 2020), GraphSIM (Yang et al., 2020), and PointSSIM (Alexiou and Ebrahimi, 2020).
  • The NR-PCQA methods include BRISQUE (Mittal et al., 2012a), NIQE (Mittal et al., 2012b), IL-NIQE (Zhang et al., 2015), IT-PCQA (Yang et al., 2022), ResSCNN (Liu et al., 2022), PQA-net (Liu et al., 2021), 3D-NSS (Zhang et al., 2022), GMS-3DQA (Zhang et al., 2023) and MM-PCQA (Zhang et al., 2023).

### v2

@@ Abstract.
Although large multi-modality models (LMMs) have seen extensive exploration and application in various quality assessment studies, their integration into Point Cloud Quality Assessment (PCQA) remains unexplored. Given LMMs’ exceptional performance and robustness in low-level vision and quality assessment tasks, this study aims to investigate the feasibility of imparting PCQA knowledge to LMMs through text supervision. To achieve this, we transform quality labels into textual descriptions during the fine-tuning phase, enabling LMMs to derive quality rating logits from 2D projections of point clouds. To compensate for the loss of perception in the 3D domain, structural features are extracted as well. These quality logits and structural features are then combined and regressed into quality scores. Our experimental results affirm the effectiveness of our approach, showcasing a novel integration of LMMs into PCQA that enhances model understanding and assessment accuracy. We hope our contributions can inspire subsequent investigations into the fusion of LMMs with PCQA, fostering advancements in 3D visual quality analysis and beyond. The code is available at https://github.com/zzc-1998/LMM-PCQA.
@@
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3680946isbn: 979-8-4007-0686-8/24/10ccs: Human-centered computing Visualization design and evaluation methodsccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Point clouds are increasingly used across diverse real-world scenarios, including virtual/augmented reality (Park et al., 2008; Liu et al., 2020; Gu et al., 2019), autonomous vehicles (Cui et al., 2021), and video post-production (Mekuria et al., 2016a). This surge is attributed to their adeptness in three-dimensional representation. Consequently, significant research efforts have been channeled towards enhancing high-level areas like point cloud classification (Grilli et al., 2017; Ku et al., 2018; Wang and Jia, 2019; Vora et al., 2020; Xie et al., 2020; Yoo et al., 2020; Chen et al., 2020), detection (Cui et al., 2021), and segmentation (Cheng et al., 2021; Liu et al., 2022a). Meanwhile, as a key component for ensuring point cloud quality, point cloud quality assessment (PCQA) has seen comparable advancements as well during the last decade. PCQA’s objective is to appraise point clouds’ visual quality, a pivotal factor for refining simplification and compression strategies in practical applications (Fan et al., 2022a), and to elevate the Quality of Experience (QoE) for end-users. Generally, PCQA methods are divided into three categories, depending on their reliance on reference point clouds: Full-Reference PCQA (FR-PCQA), Reduced-Reference PCQA (RR-PCQA), and No-Reference PCQA (NR-PCQA) respective. The availability of pristine reference point clouds is often limited in real-world applications, emphasizing the demand for NR-PCQA approaches, which is why our research primarily focuses on NR-PCQA.
@@
Secondly, to address the potential insensitivity to geometric distortions (e.g., compression, downsampling) when only analyzing projections, we propose the extraction of multi-scale structural features. This approach enhances the LMM-PCQA’s holistic comprehension of point cloud visual quality. The point clouds are converted into quality-aware structural domains, a technique validated for effective quality feature extraction in prior research (Alexiou et al., 2021; Zhou et al., 2023; Zhang et al., 2022b). We modify the scale parameters in the k-nearest neighbors (k-NN) algorithm to offer a multi-scale perspective, aligning with the human vision system’s perception mechanism. Subsequently, key statistical parameters are used to quantify structural distortions within these domains. Finally, we combine the LMM’s evaluative results and the structural features, utilizing support vector regression (SVR) to derive the quality values. The experimental outcomes affirm that our LMM-PCQA model is on par with, or surpasses, current leading PCQA methods.
@@ 2.1. LMM for Quality Assessment
Recent studies are exploring LMMs in visual quality assessment. X-IQE (Chen, 2023) uses them to evaluate text-to-image generation via a chain of thoughts strategy. Q-Bench (Wu et al., 2024a) introduces a binary softmax method for LMMs to produce quantifiable scores by assessing tokens good and poor. Q-Instruct (Wu et al., 2024b) improves this by fine-tuning LMMs through text-based questioning for specific visual queries. Q-Align (Wu et al., 2023) adopts human-like evaluation mechanisms, enhancing visual quality scoring.
@@ 2.2. PCQA Development
In the early stages of PCQA, the MPEG group developed Full-Reference PCQA (FR-PCQA) methods like p2point (Mekuria et al., 2016b) and p2plane (Tian et al., 2017) for point cloud quality assessment, and later introduced a color-based PSNR-yuv method (Torlig et al., 2018). Despite challenges in addressing complex distortions, advanced FR-PCQA metrics like PCQM (Meynet et al., 2020), GraphSIM (Yang et al., 2020b), and PointSSIM (Alexiou and Ebrahimi, 2020) were developed to integrate structural features for enhanced performance.
Building on progress in no-reference image and video quality assessment (NR-I/VQA) (Zhang et al., 2015a; Liu et al., 2015; Gu et al., 2017; Zhang and Liu, 2022; Zhang et al., 2022a, 2021a), various NR-PCQA methods have evolved. Techniques include using CNNs for quality regression with handcrafted features (Chetouani et al., 2021), multi-view projections for feature extraction (Liu et al., 2021b; Zhang et al., 2022b), and direct prediction with sparse CNNs (Liu et al., 2022c). Innovative approaches also involve converting point clouds to videos for VQA (Fan et al., 2022b; Zhang et al., 2022c), employing structure-guided resampling (Zhou et al., 2022), integrating multi-projections for simpler feature extraction (Zhang et al., 2023b), adapting image quality assessment methods for point cloud rendering (Yang et al., 2022), and using non-local geometry and color gradient models for quality estimation (Wang et al., 2023). Recent frameworks like MM-PCQA (Zhang et al., 2023a) and pmBQA (Xie et al., 2023) leverage multi-modal learning for improved PCQA. Emerging research explores text modality’s potential in quality assessment learning.
@@ 3.1.1. Projection Acquisition
Consider a colored point cloud $\mathbf{P}={(p_{i},c_{i})}_{i=1}^{N}$, where $p_{i}\in\mathbb{R}^{1\times 3}$ represents the single point consisting of geometry coordinates, $c_{i}\in\mathbb{R}^{1\times 3}$ represents the RGB color attributes, and $N$ denotes the total count of points. We then adopt the conventional cube-like viewpoints configuration, which is inspired by the projections approach discussed in (Torlig et al., 2018; Alexiou et al., 2019). As illustrated in Fig. 2, six orthogonal viewpoints are utilized, each mapping to one of the cube’s six faces for generating the projections. For a point cloud $\mathbf{P}$, the rendering process is derived as:
@@ 3.1.2. How to inject PCQA knowledge into LMMs ?
LMM has shown competitive performance in 2D image/video quality assessment (Wu et al., 2023, 2024a, 2024b), therefore it is feasible to apply LMM for PCQA tasks by taking the point cloud as a sequence of projections. Then it is vital to solve the core problem of PCQA Knowledge Injection. Following the common approaches of training LMMs (Ye et al., 2023a; Liu et al., 2023), it is natural to come up with the solution of instructing the LMM with question-answer pairs regarding PCQA issues. Thus we carry out the specific prompt structure as follows:
@@ 3.2. Structual Feature Extraction
The inadequacy of single-modality information from projections for point cloud quality evaluation has been established (Zhang et al., 2023a; Xie et al., 2023). Therefore, our approach enhances the accuracy of LMM evaluation results (projections only) by integrating geometric structural features, aiming for a more detailed and accurate assessment.
@@ 3.2.1. Structural Domain
where $Lin(p_{i})$ and $Pla(p_{i})$ represent the linearity and planarity values for point $p_{i}$. The chosen structural domains (linearity, planarity) have been demonstrated to exhibit a strong correlation with geometric visual losses, such as compression and downsampling, and have been extensively utilized in numerous PCQA tasks (Alexiou et al., 2021; Zhou et al., 2023; Zhang et al., 2022b).
@@ 3.2.2. Multi-scale Perception
The multi-scale nature of point cloud visual perception has been noted in the literature (Zhang et al., 2021b). To account for this, we compute structural domains across various scales by varying the scale parameter $k$ in the $\boldsymbol{\rm{KNN}}(\cdot)$ process. This approach allows us to derive the multi-scale structural domains as follows:
@@ 4.1. Validation Databases
To assess the efficacy of the proposed method, we employ the SJTU-PCQA database (Yang et al., 2020a), the Waterloo point cloud assessment database (WPC)(Liu et al., 2022b), and the WPC2.0 database (Liu et al., 2021a) for validation. The SJTU-PCQA database contains 9 reference point clouds, subjected to seven distortion types (compression, color noise, geometric shift, down-sampling, and three mixed distortions) at six levels, yielding a total of 378 ($9\times 7\times 6$) distorted point clouds. The WPC database includes 20 reference point clouds, each modified by four distortions (down-sampling, Gaussian white noise, Geometry-based Point Cloud Compression (G-PCC), and Video-based Point Cloud Compression (V-PCC)), resulting in 740 ($20\times 37$) distorted point clouds. Meanwhile, the WPC2.0 database features 16 reference point clouds, each undergoing 25 different V-PCC degradation settings, leading to 400 ($16\times 25$) distorted point clouds.
@@ 4.2.2. Validation Strategy
Following the methodologies in (Fan et al., 2022b; Zhang et al., 2023a; Chai et al., 2024), we utilize a k-fold cross-validation approach in our experiments to ensure a dependable performance evaluation of our proposed method. The SJTU-PCQA, WPC, and WPC2.0 databases consist of 9, 20, and 16 point cloud groups, respectively, leading us to adopt 9-fold, 5-fold, and 4-fold cross-validation for these databases to achieve an approximate 8:2 train-test split. The average of the performance metrics is considered the definitive result. It is crucial to note that the training and testing sets are mutually exclusive to prevent content overlap. For FR-PCQA methods, which do not require training, we evaluate them using the same test sets and report the average performance.
@@ 4.3. Competitors
  • The FR-PCQA methods include MSE-p2point (MSE-p2po) (Mekuria et al., 2016b), Hausdorff-p2point (HD-p2po) (Mekuria et al., 2016b), MSE-p2plane (MSE-p2pl) (Tian et al., 2017), Hausdorff-p2plane (HD-p2pl) (Tian et al., 2017), PSNR-yuv (Torlig et al., 2018), PCQM (Meynet et al., 2020), GraphSIM (Yang et al., 2020b), and PointSSIM (Alexiou and Ebrahimi, 2020).
  • The NR-PCQA methods include BRISQUE (Mittal et al., 2012a), NIQE (Mittal et al., 2012b), IL-NIQE (Zhang et al., 2015b), IT-PCQA (Yang et al., 2022), ResSCNN (Liu et al., 2022c), PQA-net (Liu et al., 2021b), 3D-NSS (Zhang et al., 2022b), GMS-3DQA (Zhang et al., 2023b) and MM-PCQA (Zhang et al., 2023a).
@@ 5. Conclusion
Acknowledgements.
This work was supported in part by the National Natural Science Foundation of China (623B2073, 62101326, 62225112, 62301316), the Foundation of Key Laboratory of Media Audio & Video (Communication University of China), Ministry of Education, China (JGKFKT2302), the China Postdoctoral Science Foundation under Grants 2023TQ0212 and 2023M742298, and the Postdoctoral Fellowship Program of CPSF under Grant GZC20231618.